OpenStudy (anonymous):

Which of the following describes the transformation of the graph y = x^2 in graphing y = -x^2 - 5? 1.) reflect over the x-axis and shift down 5 2.) reflect over the y-axis and shift down 5 3.) reflect over the x-axis and shift left 5

OpenStudy (anonymous):

Factor the negative out, for starters.

OpenStudy (anonymous):

Don't bother. Start with y = x^2 Then add the negative y = -x^2 That flips it over the x-axis Then - 5 y = -x^2 - 5 That moves it down 5 units

OpenStudy (anonymous):

Thank you!:) @BangkokGarrett